# MuPT: A Generative Symbolic Music Pretrained Transformer

### v2

@@ 3.4 Scaling Law
The Chinchilla Law, proposed by DeepMind, is a scaling law that provides insights into the training of large language models (LLMs). Our experiments reveal that the Chinchilla Law (Hoffmann et al., 2022) provides a good fit for general cases, where moderate models were trained with a moderate amount of data. In this section, we will list several improvements to Chinchilla Law for symbolic music scaling principles on limited training data.
@@ 3.4.2 Symbolic Music Scaling (SMS) Law
To enhance the predictive accuracy of the Data-Constrained law (Muennighoff et al., 2024) for continuous domains, we extend the original discrete formulation 2 to accommodate continuous variables, allowing for a more nuanced understanding of data constraints in varied contexts. For an in-depth discussion on the derivation and implications of this continuous formulation, please refer to Appendix A.2.
$L(N,D,U_{D})=\frac{A}{N^{\alpha}}+\frac{B}{{D^{\prime\prime}}^{\beta}}+E$ (2)
$D^{\prime\prime}=\frac{1-k^{D/U_{D}}}{1-k}U_{D}.$ (3)
@@
To address the model’s limitations in accurately capturing performance metrics for smaller data sizes, we introduce an additional term, as delineated in Equation 4. This modification aims to refine the model’s fidelity, particularly in scenarios characterized by limited data availability. Further details on this modification can be found in the Appendix A.3.1.
$L(N,D,U_{D})=\frac{d}{N^{\alpha}\cdot D^{\prime\prime\beta}}+\frac{A}{N^{\alpha}}+\frac{B}{D^{\prime\prime\beta}}+E.$ (4)
@@ 4.2.2 SMS Law are the Best on the Training Set
Table 3: Comparison parametric fitting performance of different scaling laws.
@@ 4.2.3 Scaled-up Performance Followed SMS Law
In our SMS Law experimentation under a computational budget of $2\times 10^{20}$ FLOPs, we initially aim to train a 2.10B (or 1.98B) parameter model across 2.82 epochs on the whole 33.6B dataset per epoch, achieving a loss of 0.5279 (or 0.5280). Engineering constraints necessitated a slight scale-down to a 1.97 billion parameter model, which, intriguingly, showed a minimal loss increase to 0.529 around 2.5 epochs. Contrary to the predictions of SMS Law, the Chinchilla Law suggests optimal performance for a 990M parameter model over 6.1 epochs. Pushing boundaries, we continuously train the 1.07B parameter model and observe overfitting returns beyond 3 epochs, validating the SMS Law’s advantages in this context. Further, we train a 4.23B parameter model that underscored the SMS Law’s predictive accuracy regarding overfitting risks, affirming its value as a strategic guide in scaling up models effectively within fixed computational constraints, beneficial for efficient model scaling decisions.
@@ 4.3.3 Subjective evaluation
Model A Model B Wins (A/B) p-value Human works MuPT 81/69 0.4237 MMT 109/41 $4.2249\times 10^{-6}$ GPT-4 119/31 $6.6315\times 10^{-9}$ Random 138/12 $4.4648\times 10^{-17}$ MuPT MMT 110/40 $4.2249\times 10^{-6}$ GPT-4 115/35 $6.6641\times 10^{-8}$ Random 131/19 $1.3618\times 10^{-13}$ MMT GPT-4 95/55 0.0093 Random 103/47 0.0001 GPT-4 Random 106/44 $2.6691\times 10^{-5}$
@@ A.1.2 Data-Constrained Law
A.2 Continuous Version of Data-Constrained Law

### v3

@@ 3.4 Scaling Law
The Chinchilla Law, proposed by DeepMind, is a scaling law that provides insights into the training of large language models (LLMs). Our experiments reveal that the Chinchilla Law (Hoffmann et al., 2022) provides a good fit for general cases, where moderate models were trained with a moderate amount of data. In this section, we will list two improvements to Chinchilla Law for symbolic music scaling principles on limited training data.
@@ 3.4.2 Symbolic Music Scaling (SMS) Law
Figure 3: Chinchilla Law prediction and loss survey in the setting with 2.1B unique tokens.
Figure 3 demonstrates the Chinchilla prediction in yellow lines and the observed loss in blue. We can tell that the Chinchilla law does not provide good results when the data volume $D$ is small when the model just begins the pre-training stage, and when $D$ is large where repeated data provides overfitting. We proposed several modifications to address these problems.
@@
The data volume $D^{\prime}$ for Data-Constrained Law Muennighoff et al. (2024) at $n$ epoch is less then $D=n\times U_{D}$. We use its approximation of Equation 18 in Appendix A of their paper instead of the standard $D^{\prime}$ for better prediction and simpler formulas. For more information please refer to Equation 11 in Appendix A.2. We denote the modified data volume as $D^{\prime\prime}$.
@@
We can observe that when that model parameter is small (e.g. $N=190M$), the Chinchilla underestimates the loss value and overestimates when the model size is large (e.g. $N=1072M$). This suggests that the coefficient $B$ in the Chinchilla formula $L=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}+E$ shall be relevant to $N$ instead of a constant.
$L(N,D^{\prime\prime})=\frac{d}{N^{\alpha}\cdot D^{\prime\prime\beta}}+\frac{A}{N^{\alpha}}+\frac{B}{D^{\prime\prime\beta}}+E.$ (2)
To address the model’s limitations in accurately capturing performance metrics for smaller data sizes, we introduce an additional term, as delineated in Equation 2. This modification aims to refine the model’s fidelity, particularly in scenarios characterized by limited data availability. Further details on this modification can be found in the Appendix A.3.1. After that, we proposed another term to predict the early stop points and overfited loss curve.
@@ 4.2.2 SMS Law are the Best on the Training Set
Table 3: Comparison of parametric fitting performance of different scaling laws.
@@ 4.2.3 Scaled-up Performance Followed SMS Law
In our SMS Law experimentation under a computational budget of $1.2\times 10^{21}$ FLOPs, we initially aim to train a 2.10B (or 1.98B) parameter model across 2.82 epochs on the whole 33.6B dataset per epoch, achieving a loss of 0.5279 (or 0.5280). Engineering constraints necessitated a slight scale-down to a 1.97 billion parameter model, which, intriguingly, showed a minimal loss increase to 0.529 around 2.5 epochs. Contrary to the predictions of SMS Law, the Chinchilla Law suggests optimal performance for a 990M parameter model over 6.1 epochs. Pushing boundaries, we continuously train the 1.07B parameter model and observe overfitting returns beyond 3 epochs, validating the SMS Law’s advantages in this context. Further, we train a 4.23B parameter model that underscored the SMS Law’s predictive accuracy regarding overfitting risks, affirming its value as a strategic guide in scaling up models effectively within fixed computational constraints, beneficial for efficient model scaling decisions.
@@ 4.3.3 Subjective evaluation
Model A Model B Wins (A/B) p-value
Human works MuPT 81/69 0.4237
MMT 109/41 $4.2249\times 10^{-6}$
GPT-4 119/31 $6.6315\times 10^{-9}$
Random 138/12 $4.4648\times 10^{-17}$
MuPT MMT 110/40 $4.2249\times 10^{-6}$
GPT-4 115/35 $6.6641\times 10^{-8}$
Random 131/19 $1.3618\times 10^{-13}$
MMT GPT-4 95/55 0.0093
Random 103/47 0.0001
GPT-4 Random 106/44 $2.6691\times 10^{-5}$
@@ A.1.2 Data-Constrained Law
A.2 Continuous Adaptation of the Data-Constrained Law.
To enhance the predictive accuracy of the Data-Constrained law (Muennighoff et al., 2024) for continuous domains, we extend the original discrete formulation 11 to accommodate continuous variables, allowing for a more nuanced understanding of data constraints in varied contexts. For an in-depth discussion on the derivation and implications of this continuous formulation, please refer to Appendix A.2.
$L(N,D,U_{D})=\frac{A}{N^{\alpha}}+\frac{B}{{D^{\prime\prime}}^{\beta}}+E$ (11)
$D^{\prime\prime}=\frac{1-k^{D/U_{D}}}{1-k}U_{D}.$ (12)